OpenStudy (anonymous):

WILL MEDAL AND FAN FOR HELP evaluate the expression. (6-2)[(2+3)+4] A. 9 B. 13 C. 24 D. 36

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what do you have so far?

OpenStudy (anonymous):

nothing i do not know how to do this one

jimthompson5910 (jim_thompson5910):

you would start evaluating the stuff in the parenthesis

jimthompson5910 (jim_thompson5910):

(6-2) = ??

OpenStudy (anonymous):

4

jimthompson5910 (jim_thompson5910):

(3+2) = ??

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

oops (2+3) = ??

jimthompson5910 (jim_thompson5910):

yeah either way, it's 5

jimthompson5910 (jim_thompson5910):

(6-2)[(2+3)+4] becomes 4*[5+4]

jimthompson5910 (jim_thompson5910):

then you have 5+4 in its own brackets (basically the same as parenthesis)

jimthompson5910 (jim_thompson5910):

so you have to do that next

OpenStudy (anonymous):

so (5+4)

jimthompson5910 (jim_thompson5910):

yeah

jimthompson5910 (jim_thompson5910):

which equals what

OpenStudy (anonymous):

9

jimthompson5910 (jim_thompson5910):

we now have 4*9

OpenStudy (anonymous):

so 36 is the answer?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thanks so much i will message you if i have anymore questions

OpenStudy (anonymous):

i made a 100 on the test